OpenStudy (anonymous):

How much H3PO4 is needed to neutralize 1 liter of a 0.75M solution of NaOH?

OpenStudy (rushwr):

1st write down the balanced equation for the neutralization reaction !

OpenStudy (rushwr):

Are u gonna reply or not ???

OpenStudy (rushwr):

\[H _{3}PO _{4} + 3NaOH \rightarrow Na _{3}PO _{4} + 3H _{2}O\]

OpenStudy (rushwr):

Now we can see that the molar ratio of H3PO4 and NaOH is 1:3 Meaning : to neutralize 3 moles of NaOH we need 1mole of H3PO4 no. of moles (n) = Concentration (C) * volume (V) 1st we have to find the moles of NaOH \[n = 1 * 0.75 = 0.75 moles\] If NaOH is 0.75 moles Then H3PO4 moles = 0.75/3 = 0.25 moles So the amount of H3PO4 moles = 0.25 moles